OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. 3/3 + i

OpenStudy (anonymous):

\[\frac{3}{3+i}*\frac{3-i}{3-i}\]\[\frac{9-3i}{9-i^2}\]\[i^2=-1\]\[\frac{9-3i}{9-(-1)}=\frac{9-3i}{10}\]\[\frac{9}{10}-\frac{3i}{10}\]